Exercise 4-1 (Video) Saddle Inc. has two types of handbags: standard and custom. The costing and wants to see how the results would differ if this companys operations controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the Standard Custom Direct labor costs Machine hours Setup hours $50,000 100,000 1,200 420 1,200 90 Total estimated overhead costs are $299,100. Overhead cost allocated to the machining activity cost pool is $192,000, and $107,100 is allocated to the machine setup activity cost pool Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e9. 12.25.) Predetermined overhead rate % of direct labor cost Compute the overhead rates using the activity-based costing approach. Machining per machine hour Machine setup per setup hour
Compute the overhead rates using the activity-based costing approach Machining Machine setup per machine hour per setup hour LINK TO TEXT LINK TO TEXT VIDEO SIMILAR EXERCISE Determine the difference in allocation between the two approaches Traditional costing Standard Custom Activity-based costing Standard Custom Click If you would like to Show Work for this question: Open Show.Work LINK TO TEXYLN
Answer #1

1. Computation of the overhead rate using the traditional ( plantwide ) approach.

Direct labour cost = ($ 100000 + $ 50000) = $150000

Estimated overhead costs = $ 299100

Predetermined Overhead Rate = ( Estimated overhead costs / Direct labour cost )

= ( $ 299100 / $ 150000 ) = 199.40% ( Answer)

Predetermined Overhead Rate 199.40%  of direct labour cost.

2. Computation of the overhead rate using the activity based costing approach.

Machining = [ Machining cost pool / ( standard machine hours + custom machine hours ) ]

= [ $ 192000 / ( 1200 + 1200 ) ]

= ( $ 192000 / 2400 ) = $ 80 per machine hour

Machine set up= [ machine set up cost pool / ( standard set up hours + custom set up hours) ]

= [ $ 107100 / ( 90 + 420 )] = ( $ 107100 / 510 ) = $ 210 per set up hour

Answer:

Machining $​​​​​​80 per machine hour.

Machine set up $ 210 per set up hour.

3. Computation of the difference in allocation of the two approaches.

Traditional costing

Standard = ( $50000 * 199.40% ) = $ 99700

Custom = ( $100000* 199.40% ) = $ 199400

Activity based costing

Standard = ( 1200 *$ 80 ) + ( 90 * $210 )

= $ 96000 + $ 18900 = $ 114900

Custom = ( 1200 * $ 80 ) + ( 420 * 210 )

= $ 96000 + $ 88200 = $ 184200